- Current Alert
- Resolved Alert
- Future Alert
The most recent ITS Alerts are listed below and are also published via the ITS Alerts RSS feed. In addition to the alerts below, a continuously updated list of the latest virus threats is provided on the right side of this page.
Last updated on July 29, 2015 at 02:54:36 PM
Thursday morning July 30th from 5:00AM to 7:00AM, ITS/TNS will be performing maintenance working on the session border controllers that support voice services to Mitel/ I2 users. During this time,...read more
Last updated on July 29, 2015 at 01:07:46 PM
On Thursday, July 30, 2015 there will be a brief outage from 4:00 — 4:30 a.m. (EST) for TrackITS -- "TrackITS Application"
Last updated on July 29, 2015 at 10:54:26 AM
There will be a brief interruption of Electronic Theses and Dissertations (eTD) and Honor Theses (eHT), along with ScholarSphere during the 5-7 a.m. (EDT) maintenance window on Thursday July 30...read more
Last updated on July 29, 2015 at 07:45:43 AM
On Thursday July 30 during the maintenance window (5:00 a.m — 7:00 a.m.), ITS will be making configuration changes to the Willard hub site wireless architecture. During the maintenance window,...read more
Last updated on July 27, 2015 at 08:04:17 PM
The problem server was identified and removed from service. Password changes should be working normally now. ITS staff investigate why the problem occurred.
Last updated on July 27, 2015 at 10:28:01 AM
The service has been restored. The front end web servers and XSan storage systems were reset.
Last updated on July 27, 2015 at 08:36:39 AM
We are currently experiencing problems logging in to eDDS and any of the Imaging services. Staff are working to resolve the issue as soon as possible.
Last updated on July 26, 2015 at 08:53:48 AM
An issue impacting Penn State’s network has been brought to the attention of the IT Service Desk. The issue caused a network access outage for multiple Penn State systems. Access...read more
Last updated on July 23, 2015 at 07:14:08 PM
By 7:12 p.m. all password changes were synchronized to all servers.
Last updated on July 23, 2015 at 02:51:33 PM
At 2:37pm systems started coming back online after it was discovered that a network hardware failure occurred. ITS is investigating the cause of this hardware failure.